# Local limits of random spanning trees in random environment

### v2

@@ 1. Introduction
Theorem 1.3 shows that there is a sharp cut-off for the local limit when $\beta=n^{\gamma}$ and $\gamma$ crosses the critical value $\gamma_{c}=1$. This stands in contrast to [10, Conjecture 1.3], where the diameter is conjectured to smoothly interpolate between $n^{1/2}$ and $n^{1/3}$ when $\gamma\in[1,4/3]$. It would be interesting to study what the local limit (if it exists) should be when $\beta$ is of order $n$. Both the proofs of Theorem 1.1 and Theorem 1.3 will make use of the fact that we are able to give sharp bounds on the effective resistance (see Definition 2.1).
@@ 2.1. Electric Networks and UST
Theorem 2.2 (Kirchhoff’s formula [8, Section 4.2]).
@@
$\boldsymbol{\mathrm{P}}(e_{1},\ldots,e_{k}\in{\mathcal{T}})=det\big{(}Y(e_{i},e_{j})_{1\leq i,j\leq k}\big{)},$ (2.4)
@@ 2.2. Coupling to a random graph
$\mathbb{P}(G_{n,p}\textrm{ is not connected}\,)\leq 4e^{-\kappa(n)}$
@@ 3. Low Disorder
In this section, we prove Theorem 1.1 and Theorem 1.3 for $\beta=\beta(n)\ll n/\log n$. We first prove lower bounds on the effective resistance using concentration arguments in Section 3.1. Then we prove matching upper bounds in Section 3.2 by constructing a good flow on the weighted graph. The concentration of the effective resistance will then allow us to conclude Theorem 1.1 and, using the method of tree moments developed in [4], we may also prove Theorem 1.3.
@@ 3.1. Lower Bounds on the Effective Resistance
Applying the Berstein inequality as in [10, Section 3.1], we obtain that for $S_{m}$ a sum of $m$ edge weights, $\beta\geq 1$ and any $0<\delta\leq 1$ we have
@@
From Rayleigh’s monotonicity principle (see [8, Section 2.4]), it follows that the effective resistance on $G$ is at least the effective resistance on the graph $G^{\prime}$ where every edge between two vertices $x,y\not\in\{u,v\}$ is contracted (in particular if $|e\cap f|=0$ the edge $f$ is also contracted). On $G^{\prime}$ the series and parallel law (see [8, Section 2.3]) give that
@@ 3.2. Upper Bounds on the Effective Resistance
We may now easily prove Theorem 1.1 for the low disorder regime of $\beta\ll n/\log n$.
@@ 3.3. Transfer-impedance
$|Y^{0}(e,f)|=\begin{cases}\frac{2}{n},&\text{if }|e\cap f|=2,\\ \frac{1}{n},&\text{if }|e\cap f|=1,\\ 0,&\text{if }|e\cap f|=0.\end{cases}$
@@
$\boldsymbol{\mathrm{P}}(e_{1},\ldots,e_{k}\in{\mathcal{T}})=(1+o(1))\prod_{i=1}^{k}\frac{\textup{{w}}(e_{i})}{\mu}\boldsymbol{\mathrm{P}}^{0}(e_{1},\ldots,e_{k}\in{\mathcal{T}}).$
@@ 4. High disorder
In this section, we will first prove the edge overlap result in Theorem 1.1 for $\beta\gg n(\log n)^{2}$. Then we will adapt these techniques to show the convergence of the local limit. The key argument is that, with high probability, edges in the MST with weights outside a window of width $2\epsilon/n$ (for some $\epsilon=o(1)$ depending on $\beta$) must also appear in the RSTRE. We remark that here high disorder refers to $\beta\gg n$ and not to the high disorder regime $\beta\gg n^{4/3}$ as in [10].
@@ 4.1. Overlap
For the moment restrict to the events $\mathcal{M}_{e}$, $\mathcal{S}_{e}$ and $\mathcal{G}_{e}$. Let $S(u)$ and $S(v)$ be the connected components of the random graph process at $p^{-}$, and w.l.o.g. assume that $|S(u)|\leq|S(v)|$. Applying Kirchhoff’s formula together with the Nash-Williams inequality (see [8, Section 2.5]) using the cutset $E(S(u),S(u)^{c})$ gives that
@@ 4.2. Local limit
We expect the high disorder local limit result to also hold when $\lambda=\lambda(n)$ is some large fixed constant instead of diverging to infinity. In (4.18) we used that, with high probability, the neighborhood $B_{M}(1,r)$ grows in size at most as fast as $60^{r}(\log n)^{r}$, however, as shown in [1], the typical size of a neighborhood is of order $r^{3}$. We are not aware of any better tail bounds, say $|B_{M}(1,r)|=O(\log n)$ for any fixed $r\geq 1$, that hold with large enough probability to make (4.18) work.

### v3

@@ 1. Introduction
Theorem 1.3 shows that there is a sharp cut-off for the local limit when $\beta=n^{\gamma}$ and $\gamma$ crosses the critical value $\gamma_{c}=1$. This stands in contrast to [13, Conjecture 1.3], where the diameter is conjectured to smoothly interpolate between $n^{1/2}$ and $n^{1/3}$ when $\gamma\in[1,4/3]$. It would be interesting to study what the local limit (if it exists) should be when $\beta$ is of order $n$. Both the proofs of Theorem 1.1 and Theorem 1.3 will make use of the fact that we are able to give sharp bounds on the effective resistance (see Definition 2.1).
Very shortly after we posted this paper, Kúsz [8] uploaded a paper studying the same model as ours. In particular, using the Aldous-Broder algorithm, they manage to (independently) prove Theorem 1.1 of [13]. Furthermore, they study yet another observable called the total length, which in our notation is defined as
$L({\mathcal{T}}):=\sum_{e\in{\mathcal{T}}}\omega_{e},$
and is equivalent to the Hamiltonian $H({\mathcal{T}},\omega)$ as in [13]. In the following, we state an improvement to the low disorder part of [8, Theorem 1.8].
Theorem 1.4.
Let ${\mathcal{T}}$ be the RSTRE on the complete graph with $n$ vertices and random environment distributed as uniform random variables on $[0,1]$. If $0<\beta=\beta(n)\ll n/\log n$, then with high probability (and in $\mathbb{P}$-expectation)
$\boldsymbol{\mathrm{E}}^{\omega}_{n,\beta}\big{[}L({\mathcal{T}})\big{]}=(1+o(1))\frac{n}{\beta}\cdot\frac{1-\beta e^{-\beta}-e^{-\beta}}{1-e^{-\beta}}.$ (1.7)
If on the other hand $\beta\gg n(\log n)^{5}$, then with high probability (and in $\mathbb{P}$-expectation)
$\boldsymbol{\mathrm{E}}^{\omega}_{n,\beta}\big{[}L({\mathcal{T}})\big{]}=(1+o(1))\zeta(3),$ (1.8)
where $\zeta(3)=\sum_{k=1}^{\infty}k^{-3}=1.202\dots$ is Apéry’s constant.
The constant $\zeta(3)$ is precisely the limiting total length of the MST on the complete graph, as determined in [5]. Theorem 1.4 will be a consequence of the tools we develop for the edge overlap, see the end of Sections 3.2 and 4.1 for a (sketch) of the proofs. We remark that in [8, Theorem 1.8], it is shown that the equality (1.8) in the high disorder holds under the milder condition $\beta\gg n\log n$. See also the PhD thesis [11] of the author for yet another alternative simple proof of this fact. Additionally, in a newer version of [8], it is proven that the RSTRE local limit agrees with that of the MST local limit under the milder condition $\beta\gg n\log n$, giving an improvement to the high disorder regime of Theorem 1.3.
@@ 2.1. Electric Networks and UST
Theorem 2.2 (Kirchhoff’s formula [10, Section 4.2]).
@@
$\boldsymbol{\mathrm{P}}(e_{1},\ldots,e_{k}\in{\mathcal{T}})=\det\big{(}Y(e_{i},e_{j})_{1\leq i,j\leq k}\big{)},$ (2.4)
@@ 2.2. Coupling to a random graph
$\mathbb{P}(G_{n,p}\textrm{ is not connected}\,)\leq 4e^{-\kappa(n)}.$
@@ 3. Low Disorder
In this section, we prove Theorems 1.1, 1.3 and 1.4 for $\beta=\beta(n)\ll n/\log n$. We first prove lower bounds on the effective resistance using concentration arguments in Section 3.1. Then we prove matching upper bounds in Section 3.2 by constructing a good flow on the weighted graph. The concentration of the effective resistance will then allow us to conclude Theorem 1.1 and, using the method of tree moments developed in [4], we may also prove Theorem 1.3.
@@ 3.1. Lower Bounds on the Effective Resistance
Applying the Berstein inequality as in [13, Section 3.1], we obtain that for $S_{m}$ a sum of $m$ edge weights, $\beta\geq 1$ and any $0<\delta\leq 1$ we have
@@
From Rayleigh’s monotonicity principle (see [10, Section 2.4]), it follows that the effective resistance on $G$ is at least the effective resistance on the graph $G^{\prime}$ where every edge between two vertices $x,y\not\in\{u,v\}$ is contracted (in particular if $|e\cap f|=0$ the edge $f$ is also contracted). On $G^{\prime}$ the series and parallel law (see [10, Section 2.3]) give that
@@ 3.2. Upper Bounds on the Effective Resistance
We may now easily prove Theorem 1.1 and Theorem 1.4 for the low disorder regime of $\beta\ll n/\log n$.
@@
Proof of (1.7) in Theorem 1.4.
The expected (w.r.t. $\boldsymbol{\mathrm{E}}$) edge length satisfies
$\boldsymbol{\mathrm{E}}\big{[}L({\mathcal{T}})\big{]}=\sum_{(u,v)\in E}\omega_{(u,v)}\boldsymbol{\mathrm{P}}\big{(}(u,v)\in{\mathcal{T}}\big{)}=\sum_{(u,v)\in E}\omega_{(u,v)}e^{-\beta\omega_{(u,v)}}{\rm R}_{\rm eff}(u\leftrightarrow v).$
Therefore, by Lemma 3.1 and Lemma 3.5, we have
$\boldsymbol{\mathrm{E}}\big{[}L({\mathcal{T}})\big{]}=(1+o(1))\frac{2}{n\mu}\sum_{e\in E}\omega_{e}e^{-\beta\omega_{e}}$ (3.12)
with probability, say, at least $1-O(n^{-2})$. Let $X_{e}=\omega_{e}e^{-\beta\omega_{e}}$, then
$\mathbb{E}[X_{e}]=\frac{1-\beta e^{-\beta}-e^{-\beta}}{\beta^{2}},$
and there exists a constant $C>0$ such for $\beta\geq 1$, we have
${\rm Var}(X_{e})\leq C\frac{1}{\beta^{3}}.$
As (3.12) holds with probability at least $1-O(n^{-2})$ and the length of ${\mathcal{T}}$ is trivially between 00 and $n-1$, the equality in (1.7) holds in $\mathbb{P}$-expectation. Furthermore, using Bernstein’s inequality similarly to (3.1), with $m=n(n-1)/2$ and $m/\beta^{2}\rightarrow\infty$, one may show that right-hand side of (3.12) is, with high probability, as in (1.7). ∎
@@ 3.3. Transfer-impedance
$|Y^{0}(e,f)|=\begin{cases}\frac{2}{n}&\text{if }|e\cap f|=2,\\ \frac{1}{n}&\text{if }|e\cap f|=1,\\ 0&\text{if }|e\cap f|=0.\end{cases}$
@@
$\boldsymbol{\mathrm{P}}(e_{1},\ldots,e_{k}\in{\mathcal{T}})=(1+o(1))\prod_{i=1}^{k}\frac{\textup{{w}}(e_{i})}{\mu}\boldsymbol{\mathrm{P}}^{0}(e_{1},\ldots,e_{k}\in{\mathcal{T}})$
@@ 4. High disorder
In this section, we will first prove Theorem 1.1 and Theorem 1.4 for $\beta\gg n(\log n)^{2}$. Then we will adapt these techniques to show the convergence of the local limit. The key argument is that, with high probability, edges in the MST with weights outside a window of width $2\epsilon/n$ (for some $\epsilon=o(1)$ depending on $\beta$) must also appear in the RSTRE. We remark that here high disorder refers to $\beta\gg n$ and not to the high disorder regime $\beta\gg n^{4/3}$ as in [13].
@@ 4.1. Overlap
For the moment restrict to the events $\mathcal{M}_{e}$, $\mathcal{S}_{e}$ and $\mathcal{G}_{e}$. Let $S(u)$ and $S(v)$ be the connected components of the random graph process at $p^{-}$, and w.l.o.g. assume that $|S(u)|\leq|S(v)|$. Applying Kirchhoff’s formula together with the Nash-Williams inequality (see [10, Section 2.5]) using the cutset $E(S(u),S(u)^{c})$ gives that
@@
In the following, we briefly sketch how to obtain the equality (1.8) in the high disorder regime of Theorem 1.4. We also refer to [8, Theorem 1.8], where they only require that $\beta\gg n\log n$ instead of $\beta\gg n(\log n)^{5}$.
Proof sketch of (1.8) in Theorem 1.4.
Denote by $M$ the MST coupled to the environment as in Section 2.2. As the MST has the minimum length given the environment $\omega$, we have that
$L(M)\leq L({\mathcal{T}})\leq L(M)+|{\mathcal{T}}\setminus M|\max_{e\in{\mathcal{T}}}\omega_{e},$
where $|{\mathcal{T}}\setminus M|$ is the number of edges in ${\mathcal{T}}$ that are not in $M$. Let $\epsilon$ and $\gamma$ be as in (4.8), then the above proof of Theorem 4.1 shows that the overlap between $M$ and ${\mathcal{T}}$ satisfies
$\boldsymbol{\mathrm{E}}\big{[}|M\cap{\mathcal{T}}|\big{]}\geq\big{(}1-O(\max\{\epsilon,\gamma\})\big{)}n,$
with high probability. In particular,
$|{\mathcal{T}}\setminus M|=O(\max\{\epsilon,\gamma\})n$
with high probability. As $G_{n,p}$ is connected with high probability when $p=2\log n/n$, using Lemma 4.3 of [13], one may easily obtain that $\max_{e\in{\mathcal{T}}}\omega_{e}\leq 3\log n/n$ with high probability. Equation (1.8) now follows by letting $\epsilon,\gamma=o((\log n)^{-1})$, which is possible provided that $\beta\gg n(\log n)^{5}$. ∎
@@ 4.2. Local limit
The argument after (4.18) also shows that the neighborhoods $B_{\mathcal{T}}(u,1)$ and $B_{M}(u,1)$ agree with each other for diverging $r$, as long as $\beta$ is large enough such that
$60^{r}(\log n)^{r}\Big{(}\frac{n(\log n)^{5}}{\beta}\Big{)}^{1/3}\longrightarrow 0$
We expect that this condition can be somewhat relaxed. Indeed, for (4.18), we used the fact that, with high probability, the neighborhood $B_{M}(1,r)$ grows in size at most as fast as $60^{r}(\log n)^{r}$. However, as shown in [1], the typical size of a neighborhood is of order $r^{3}$. We are not aware of any better tail bounds, such as $|B_{M}(1,r)|=O(r^{k}\log n)$ for some $k\geq 1$, that hold with sufficiently high probability to make (4.18) work for smaller $\beta$.